1 2/3 of what number is 13
Answer by stanbon(75887)

1 2/3 of what number is 13
--------------------------------
Let the number be "x".
Equation:
[1 2/3]x = 13
(5/3)x = 13
x = (3/2)13
x = 39/2
==========================
Cheers,
Stan H.